OpenStudy (anonymous):

1.Solve S = 4v2 for v.

jhonyy9 (jhonyy9):

S=4^2 do you mean hence ?

jhonyy9 (jhonyy9):

sorry v squared

jhonyy9 (jhonyy9):

S=4v^2

OpenStudy (anonymous):

yes squared. my bad. lol

jhonyy9 (jhonyy9):

than make squareroot from both sides sqrtS =2v divide both sides by 2 sqrtS ---- = v 2

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

I don't understand, Will you show work and help me out? PLEASE :)

OpenStudy (anonymous):

\[\sqrt{s}= \sqrt{4v^2}\] \[\sqrt{s}= 2v\]

OpenStudy (anonymous):

\[\sqrt{s}/2=v\]

OpenStudy (anonymous):

ok???

OpenStudy (anonymous):

Is that showing work? LOL! I've been out of school for about 6 months and have forgotten a whole lot :'(

OpenStudy (anonymous):

its showing work .... what do u want more in it??? i m sorry if it didnt work.... but u should appreciate Duh!